# Enhancing Robustness of CLIP to Common Corruptions through Bimodal Test-Time Adaptation

### v1

@@ Abstract
Although open-vocabulary classification models like Contrastive Language Image Pretraining (CLIP) have demonstrated strong zero-shot learning capabilities, their robustness to common image corruptions remains poorly understood. Through extensive experiments, we show that zero-shot CLIP lacks robustness to common image corruptions at increasing severity levels during test-time, necessitating the adaptation of CLIP to unlabeled corrupted images using test-time adaptation (TTA). However, we found that existing TTA methods have severe limitations in adapting CLIP due to their unimodal nature. To address these limitations, we propose BAT-CLIP, a bimodal TTA method specially designed to improve CLIP’s robustness to common image corruptions. The key insight of our approach is not only to adapt the visual encoders for better image feature extraction but also to strengthen the alignment between image and text features by promoting a stronger association between the image class prototype, computed using pseudo-labels, and the corresponding text feature. We evaluate our approach on benchmark image corruption datasets and achieve state-of-the-art results in TTA for CLIP, specifically for domains involving image corruption. Particularly, with a ViT-B/16 vision backbone, we obtain mean accuracy improvements of 9.7%, 5.94%, and 5.12% for CIFAR-10C, CIFAR-100C, and ImageNet-C, respectively.
@@ 1 Introduction
Figure 1: Comparison of BAT-CLIP with other online TTA approaches using CLIP: TPT [38], and VTE [9]. a) TPT optimizes text prompts only for a single test image, making it unimodal. b) VTE considers an ensemble of prompts without model updates. Both methods consider the generation of multiple augmentations of the test image. c) BAT-CLIP is a bimodal approach, that adapts the LayerNorm parameters of the vision and text encoders, maximizing alignment between class prototypes and text features while increasing the inter-class separability of prototypes.
@@
Despite CLIP’s successes in such important applications, its robustness when faced with corrupted images remains largely underexplored. Our motivation stems from the fact that the vision perception system of humans exhibits a level of robustness that real-world vision systems are yet to achieve. For example, models deployed for safety-critical applications like autonomous driving [1], could face rapid distributional shifts of blurriness, pixel changes, snowy nights, or other weather conditions [35]. For instance, our findings on the zero-shot performance of CLIP with a ResNet-101 [14] vision backbone reveals that the accuracy on the test set of CIFAR100 [17] with Gaussian noise of severity level 5, plummets to 10.79% from 49% on the clean set. Similar trends are observed with ViT-B/16, -B/32, and -L/14 [10] as backbones. Such behaviors can result in significant performance degradation of models when encountering image corruption in real-world scenarios, particularly in applications like self-driving cars.
This challenge is not unique to CLIP; it reflects a broader issue in computer vision tasks [7, 11, 12, 5], which often rely on the assumption that training and test data share the same distribution. When test distributions differ, model adaptation becomes essential to accommodate this shift and maintain accurate predictions. In response to this issue, test-time adaptation (TTA) has garnered significant interest [46, 8]. TTA aims to adapt a pre-trained model to unlabeled batches of test data from varying domains in an online fashion—updating the model for each test batch without access to the source dataset. Several TTA methods proposed in the literature have been effective in mitigating domain shifts [46, 36, 34, 28, 42, 3]. In this paper, we use the terms distributions and domains interchangeably.
@@
While there have been few new works on using CLIP for online TTA [38, 9], they come with certain limitations. For example, test-time prompt tuning (TPT) [38] tunes the text prompts on the text encoder alone and generates multiple random augmented views, for each test image. The text prompts, initialized to pre-trained values, are optimized by minimizing the marginal entropy of the confident model predictions. The prompts are reset after adaptation to each image. However, such a method is expensive and slow due to performing multiple forward passes through the vision encoder of CLIP, for each image. Also, it relies on hand-crafted prompts for initialization, making it impractical at test-time. Another recent approach, Vision-Text-Space Ensemble (VTE) [9], uses an ensemble of different prompts as input to CLIP’s text encoder while keeping both the encoders frozen. Since the vision encoder is frozen, it struggles to adapt images with severe noise effectively.
@@
While TPT effectively improves CLIP’s test generalization by dynamically tuning the text prompts, it remains primarily a unimodal approach. This limits the capacity of a multimodal model like CLIP to fully leverage its multimodal nature for adaptation. Specifically, it prevents the encoders from jointly adjusting their features, resulting in suboptimal alignment between the visual and text modalities. During TTA, there is no transfer of knowledge, via gradients, between the visual and text encoders, as the adaptation process focuses solely on one modality. For instance, when a test image includes common corruptions, the text prompts adapt, but the vision encoder features remain fixed. As a result, the learned prompts lack awareness of the test image distribution, leading to a less effective adaptation.
@@
To address these core limitations, we propose, BAT-CLIP, a bimodal adaptation approach of CLIP for online TTA, where both the visual and text encoders are jointly adapted by exploiting CLIP’s shared feature space of images and text. The overall objective is to achieve a strong alignment between image features and text features to enable more effective multimodal learning and adaptation. The adaptation procedure is two-fold: 1) Inspired by [47] for efficient fine-tuning, we adapt the norm layers i.e., LayerNorm parameters of both encoders. However, such a model update does not consider the alignment of the encoder features. So, to improve the alignment between class-specific visual and text features, we introduce a projection matching loss that maximizes the projection of the visual class prototypes with their corresponding text features. 2) To learn more discriminative visual features, we increase the cosine distance between the class prototypes, promoting a more distinct separation in the image feature space. Furthermore, our method is general-purpose, as it does not rely on hand-crafted prompt templates or their ensembles, unlike VTE. We leverage batches of test samples for TTA, rather than focusing on single-image adaptation, which makes our method fast. In the Supplementary, we show several examples of classification results, as a comparison.
Our novelty lies in performing a bimodal online test-time adaptation (TTA) of CLIP for classification tasks, to improve CLIP’s robustness. We draw a comparison of our method against TPT and VTE in Figure 1. Our main contributions are as follows:
@@
  • We start with an in-depth analysis of CLIP’s zero-shot performance across different visual backbones on common image corruptions, at test-time, for varying severity levels. While CLIP shows strong performance on clean images, its performance drops notably on corrupted images.
@@
  • To address the unimodal limitations highlighted earlier, our proposed bimodal adaptation of CLIP encoders at test-time aims to enhance alignment by maximizing the projection of class-wise prototypes onto their corresponding text features. Simultaneously, we increase the cosine distance between class prototypes to encourage the learning of more discriminative features, making the test adaptation process more flexible and robust.
@@
  • We conduct extensive experiments benchmarking our method against prior TTA approaches using CLIP, i.e., TPT and VTE, for online test-time adaptation. We also compare BAT-CLIP against four TTA baselines, adopted for CLIP. Our approach achieves state-of-the-art performance for CLIP adaptation to common corruptions.
@@ 2 Related Works
Online Test-Time Adaptation (TTA). The objective of TTA is to adapt a pre-trained model, with no access to the source data, to incoming batches of unlabelled test data of a specific domain [46, 42, 36, 34, 28, 3, 49, 6]. In this approach, the model is reset to its pre-trained state after adapting to each target domain. As the updates are performed online, TENT [46] updates the affine parameters of the normalization layers and minimizes the entropy [37] of the model predictions. RoTTA [47] proposes a robust feature normalization method, leveraging a memory bank to track test data distribution and a teacher-student framework for adaptive reweighting and updates. RPL [34] argues that self-learning during adaptation via entropy minimization and pseudo-labels is beneficial. They propose the usage of a generalized cross-entropy loss for adaptation. SAR [28] filters noisy test samples that cause a performance drop identified from the gradient space. However, none of these works utilize CLIP for TTA.
TTA using CLIP. Lately, CLIP has been finding extensive applications for TTA. TPT [38] was the first work to propose prompt tuning using CLIP at test-time in an online manner. However, this method is computationally intensive for generating multiple views, per image. Additionally, a key limitation of this approach is its reliance on using CLIP’s default hand-crafted template for prompt initialization, with resets after adaptation to each image. A similar line of work was done in VTE [9] where ensembles are created in the text and vision space without any CLIP parameter update. [40] focus on single-image TTA for out-of-distribution detection and test generalization [22, 19]. WATT [29] borrows the idea of “test-time training” i.e., over several optimization steps, it adapts CLIP’s vision encoder using multiple text prompt templates, for a batch of test images. However, this TTA setting differs from ours where we update the model only once on a single test batch. SwapPrompt [26] explores the idea of self-supervised contrastive learning for prompt learning. HisTPT [48] relies on a memory bank, to store features of recent test samples, for prompt learning. Such a setup is impractical in real-time and is expensive. To summarize — while TPT is unimodal, our approach is bimodal, involving on-the-fly joint optimization of both encoders for test adaptation, leading to a strong multimodal alignment between features. We use a generic prompt template, avoiding prompt engineering and enabling real-time deployment.
@@ 3 Zero-shot performance analysis of CLIP to common image corruptions
Vision Backbones. We evaluate the robustness with a ResNet-101 (RN101) vision backbone [14] and three Vision Transformer backbones (ViT-B/16, ViT-B/32, ViT-L/14) [10]. The models are tested on their zero-shot classification performance i.e., accuracy.
Datasets. For all the experiments in this paper, we utilize the CIFAR-10C, CIFAR-100C, and ImageNet-C [15] datasets, which are standard benchmark datasets to evaluate the robustness of classification models. Each dataset contains 15 distinct corruption types as tasks (e.g., Gaussian noise, Shot noise, …). Each corruption is applied at 5 different severity levels to the test sets of CIFAR10, CIFAR100 [17], and ImageNet [7], acting as source test sets, and allowing us to systematically evaluate the model’s performance under increasing degrees of image degradation.
@@
where $\text{sim}(\cdot)$ is the cosine similarity and $\tau$ is the softmax temperature from CLIP’s pre-training stage. The text features $\mathcal{Z}$, in zero-shot evaluation, are always pre-computed. We also draw a comparison with CLIP performance on respective source test sets and follow  [^1] https://github.com/LAION-AI/CLIP_benchmark for this implementation. Throughout this paper, the same TTA problem setup is employed.
Figure 2: Task-wise mean accuracy (%) of zero-shot CLIP across corruption severity levels. [Top]: ResNet-101 backbone. [Bottom]: ViT-B/16 backbone. Dashed lines show zero-shot CLIP performance on corresponding source test sets.
Prompt Template Backbone
CIFAR-10C
CIFAR-100C
ImageNet-C
“a photo of a ¡CLS¿.”
RN101
42.30 (-36.50) 20.62 (-25.48) 14.97 (-46.23)
ViT-B/16
61.16 (-28.84) 35.78 (-30.82) 24.51 (-42.16)
ViT-B/32
59.00 (-29.20) 31.79 (-30.51) 23.20 (-38.80)
ViT-L/14
75.84 (-19.36) 47.82 (-27.78) 39.55 (-33.95)
“a bad photo of a ¡CLS¿.”
RN101
42.73 (-36.47) 20.18 (-27.02) 12.87 (-49.43)
ViT-B/16
62.97 (-28.03) 36.37 (-30.33) 25.84 (-42.26)
ViT-B/32
60.25 (-29.75) 31.30 (-31.00) 23.58 (-39.22)
ViT-L/14
76.08 (-19.22) 48.09 (-25.41) 39.37 (-34.73)
“a blurry photo of a ¡CLS¿.”
RN101
45.32 (-34.98) 20.28 (-26.12) 15.80 (-45.70)
ViT-B/16
62.49 (-28.31) 35.17 (-30.83) 25.27 (-42.03)
ViT-B/32
57.85 (-30.65) 31.47 (-30.85) 23.40 (-38.80)
ViT-L/14
73.77 (-20.73) 48.07 (-25.73) 39.22 (-34.18)
“a noisy photo of a ¡CLS¿.”
RN101
44.56 (-34.34) 20.71 (-25.89) 15.48 (-45.72)
ViT-B/16
63.03 (-27.87) 35.14 (-30.76) 25.05 (-41.35)
ViT-B/32
59.76 (-29.24) 31.53 (-30.37) 23.11 (-38.19)
ViT-L/14
75.36 (-19.54) 47.87 (-27.63) 38.78 (-34.42)
Table 1: Mean classification accuracy (%) across all corruption tasks with various prompt templates. We report the accuracy drop (-) (%) relative to performance on the corresponding source test set.
@@ 3.1 Sensitivity of CLIP to image corruption severity
We analyze CLIP’s zero-shot performance by progressively increasing corruption severity for RN101 and ViT-B/16 vision backbones. We report mean accuracy and overall performance in Figure 2. Despite CLIP’s robust multimodal feature space, accuracy drops significantly with an increase in corruption severity, regardless of the backbone. Specifically, for CIFAR-10C with a ViT-B/16 backbone, we observe accuracy as low as 37.92%, at a severity level of 5, for Gaussian noise. Similarly, for CIFAR-100C and ImageNet-C, the mean accuracy rates are as low as 35.79% and 24.51% at a severity level of 5, respectively. Results on additional vision backbones are in the Supplementary.
@@ 3.2 Sensitivity of CLIP to prompt templates
In this analysis, we evaluate the impact of prompt engineering by providing “relevant” prompt templates to $f_{txt}$ for TTA. Each prompt adds context to help CLIP extract more relevant text features. We report the mean classification accuracy (in %) across RN101, ViT-B/16, ViT-B/32, and ViT-L/14 backbones at an image corruption severity level of 5 for all datasets, with results summarized in Table 1. The absolute accuracy drop compared to zero-shot performance on the source test set is also reported.
@@
As expected, RN101 performs worse than the ViT-based backbones, primarily due to its lack of global attention-based modeling inherent to transformers [44]. Therefore, for the remainder of the experiments in this paper, we focus on ViT-based backbones, specifically ViT-B/16 and ViT-B/32. Due to GPU memory limitations, we do not use ViT-L/14 for model adaptation.
Unsuitability of prompt template selection at test-time. Additionally, at test-time, it is impractical to perform prompt engineering or optimize prompt vectors since 1) Choosing different prompt templates for generating text features is extremely tedious and time-consuming. 2) In real-time deployment involving prompt-tuning, such prompts cannot quickly estimate the distribution of incoming test batches. TPT [38] optimizes pre-trained text prompts for each test image, turning out to be suboptimal since the prompts are optimized ignorant of the distribution of the test image.
4 BAT-CLIP
Figure 3: BAT-CLIP not only adapts the visual encoder for highly discriminative image features but also promotes a strong alignment between image and text features by adapting the text encoder too, leading to improved performance following test-time adaptation. We adapt only the LayerNorm parameters of CLIP encoders.
The comprehensive analysis in Section 3 reveals that the zero-shot vision encoder $f_{vis}$ of CLIP is very sensitive to image corruption with increasing severity. Similarly, the performance of the text encoder $f_{txt}$ is invariant to the different text prompt templates and is also impractical to tune at test-time. Therefore, for effective adaptation, both the image and text features of CLIP need to be adapted to the incoming domain of test batches. We illustrate our proposed framework, BAT-CLIP, in Figure 3.

### v2

# BATCLIP: Bimodal Online Test-Time Adaptation for CLIP

@@ Abstract
Although open-vocabulary classification models like Contrastive Language Image Pretraining (CLIP) have demonstrated strong zero-shot learning capabilities, their robustness to common image corruptions remains poorly understood. Through extensive experiments, we show that zero-shot CLIP lacks robustness to common image corruptions during test-time, necessitating the adaptation of CLIP to unlabeled corrupted images using test-time adaptation (TTA). However, we found that existing TTA methods have severe limitations in adapting CLIP due to their unimodal nature. To address these limitations, we propose BATCLIP, a bimodal online TTA method designed to improve CLIP’s robustness to common image corruptions. The key insight of our approach is not only to adapt the visual encoders for improving image features but also to strengthen the alignment between image and text features by promoting a stronger association between the image class prototype, computed using pseudo-labels, and the corresponding text feature. We evaluate our approach on benchmark image corruption datasets and achieve state-of-the-art results in online TTA for CLIP. Furthermore, we evaluate our proposed TTA approach on various domain generalization datasets to demonstrate its generalization capabilities.
@@ 1 Introduction
Figure 1: Comparison of BATCLIP with other online TTA approaches using CLIP. a) TPT [45] optimizes text prompts only for a single test image, making it unimodal. b) VTE [10] considers an ensemble of prompts without model updates. Both methods consider the generation of multiple augmentations of the test image. c) BATCLIP is a bimodal approach, that adapts the LayerNorm parameters of the vision and text encoders, maximizing alignment between class prototypes and text features while increasing the inter-class separability of prototypes.
@@
Despite CLIP’s successes in such important applications, its robustness when faced with corrupted images remains largely underexplored. Our motivation stems from the fact that the vision perception system of humans exhibits a level of robustness that real-world vision systems are yet to achieve. For example, models deployed for safety-critical applications like autonomous driving [1], could face rapid distributional shifts of blurriness, pixel changes, snowy nights, or other weather conditions [42]. In particular, our findings on the zero-shot performance of CLIP with a ResNet-101 [17] vision backbone reveals that the accuracy on the test set of CIFAR100 [22] with Gaussian noise of severity level 5, plummets to 10.79% from 49% on the clean set. Similar trends are observed with ViT-B/16, -B/32, and -L/14 [11] as backbones. The performance degradation caused by image corruption can have significant consequences in real-world scenarios, particularly in safety-critical applications like self-driving cars.
One effective approach to improving model performance under distribution shift is test-time adaptation (TTA) [53, 9], where a pre-trained model is directly adapted to unlabeled test batches, without access to the source dataset. Moreover, given the abundance of unlabeled data available in the wild, there is a growing need for online adaptation to streaming data, with a single forward pass, ensuring privacy, real-time performance [31], and preventing catastrophic forgetting [13] of source knowledge. Several online TTA methods proposed in the literature have been effective in mitigating domain shifts [53, 43, 41, 35, 48, 4]. In this paper, we use the terms distributions and domains interchangeably.
@@
While there have been few works on using CLIP for online TTA [45, 10, 21], they come with certain limitations. For example, test-time prompt tuning (TPT) [45] tunes the text prompts on the text encoder alone and generates multiple random augmented views, for each test image. The text prompts, initialized to pre-trained values, are optimized by minimizing the marginal entropy of the confident model predictions. The prompts are reset after adaptation to each image. However, such a method is expensive and slow due to performing multiple forward passes through the vision encoder of CLIP, for each image. Also, it relies on hand-crafted prompts for initialization, making it impractical at test-time. Another recent approach, Vision-Text-Space Ensemble (VTE) [10], uses an ensemble of different prompts as input to CLIP’s text encoder while keeping both the encoders frozen. Since the vision encoder is frozen, it struggles to effectively adapt to images with severe noise.
@@
While TPT effectively improves CLIP’s test generalization by dynamically tuning the text prompts, it remains primarily a unimodal approach. This limits the capacity of a multimodal model like CLIP to fully leverage its multimodal nature for adaptation. Specifically, it prevents the encoders from jointly adjusting their features, resulting in suboptimal alignment between the visual and text modalities after adaptation. For instance, when a test image includes common corruptions, the text prompts adapt, but the vision encoder features remain fixed. As a result, the learned prompts lack awareness of the test image distribution, leading to a less effective adaptation.
@@
To address these core limitations, we propose, BATCLIP, a bimodal adaptation approach of CLIP for online TTA specifically, where both the visual and text encoders are jointly adapted by exploiting CLIP’s shared feature space of images and text. The overall objective is to achieve a strong alignment between image features and text features to enable more effective multimodal learning and adaptation. The adaptation procedure is two-fold: 1) Inspired by [54, 58] for efficient fine-tuning, we only adapt the norm layers, i.e., LayerNorm parameters of both encoders. However, such a model update does not consider the alignment of the encoder features. Therefore, to improve the alignment between class-specific visual and text features, we introduce a projection matching loss that maximizes the projection of the visual class prototypes with their corresponding text features. 2) To learn more discriminative visual features, we enhance the cosine distance between class prototypes, computed using pseudo-labels, to promote a more distinct separation in the image feature space. Our method is general-purpose, with no reliance on multiple prompt templates [36, 15, 34] or their ensembles, unlike VTE. We leverage batches of test samples for TTA, rather than focusing on single-image test adaptation, which makes our method fast. In the Supplementary, we show several examples of classification results, as a comparison. Figure 1(d) shows that with a bimodal online test-time adaptation (TTA) setup for CLIP, the proposed approach achieves a significant improvement compared to TPT and VTE.
Our main contributions are as follows:
@@
  • We start with an in-depth analysis of CLIP’s zero-shot performance across different visual backbones on an established benchmark on common image corruptions [18], at test-time, for varying severity levels. We found that while CLIP shows strong performance on clean images, its performance drops notably on corrupted images.
@@
  • To address the unimodal limitations highlighted earlier, we propose BATCLIP, a bimodal online test-time adaptation method for CLIP encoders, designed to enhance alignment by maximizing the projection of class-wise prototypes onto their corresponding text features. Simultaneously, we increase the cosine distance between class prototypes to encourage learning of more discriminative features, making the test adaptation process more flexible and robust.
@@
  • We conduct extensive experiments benchmarking the proposed method against established TTA baselines and others using CLIP on CIFAR-10C, CIFAR-100C, and ImageNet-C [18]. To generalize our method beyond common corruptions, we evaluate on various domain generalization datasets [14]-OfficeHome [51], PACS [25], VLCS [12], and Terra Incognita [3]. Overall, our approach achieves state-of-the-art performance for CLIP adaptation at test-time on various domain shifts.
@@ 2 Related Works
Online Test-Time Adaptation (TTA). The objective of online TTA is to adapt a pre-trained model, with no access to the source data, to incoming batches of unlabelled test data of a specific domain [53, 48, 43, 41, 35, 4, 57, 6]. In this approach, the model is reset to its pre-trained state after adapting to each target domain. As the updates are performed online, TENT [53] updates the affine parameters of the normalization layers and minimizes the entropy [44] of the model predictions. RoTTA [54] proposes a robust feature normalization method, leveraging a memory bank to track test data distribution and a teacher-student framework for adaptive reweighting and updates. RPL [41] argues that self-learning during adaptation via entropy minimization and pseudo-labels is beneficial. They propose the usage of a generalized cross-entropy loss for adaptation. SAR [35] filters noisy test samples that cause a performance drop identified from the gradient space. However, none of these works utilize CLIP for TTA.
[…]
Another set of TTA works using CLIP [36, 15, 34] leverage information from multiple prompt templates. WATT [36] adapts CLIP’s vision encoder, weight-averaging the adapted weights from multiple prompt templates, for each test batch, over several optimization steps. Similarly, CLIPArTT [15] enhances text supervision from the support of pseudo-labels. Mishra et al. [34] pose pseudo-labeling as an Optimal Transport [7] problem by distilling knowledge from multiple text prototypes, for several iterations. Please note that the TTA setups here largely differ from our online TTA setup [53].
While promising, these approaches have key limitations. Firstly, multiple gradient descent steps on a test batch can result in large parameter shifts, leading to the loss of CLIP’s source knowledge i.e., catastrophic forgetting [13]. No regularization strategies have been proposed to restrict the loss of CLIP’s pretrained knowledge. Secondly, these methods are not favorable for real-time deployment where fast and on-the-fly test-adaptation must meet privacy and memory constraints. Thirdly, relying on multiple prompt templates, at test-time, is impractical. We discuss this in Section 3.
@@ 3 Zero-shot performance analysis of CLIP to common image corruptions
Vision Backbones. We evaluate the robustness with a ResNet-101 (RN101) vision backbone [17] and three ViT backbones (ViT-B/16, ViT-B/32, and ViT-L/14) [11]. The models are tested on their zero-shot classification performance.
Datasets. For all the experiments in this paper, we utilize the CIFAR-10C, CIFAR-100C, and ImageNet-C [18] datasets, which are standard benchmark datasets to evaluate the robustness of vision models [18]. Each dataset contains 15 distinct corruption types as tasks (e.g., Gaussian noise, Shot noise, …). Each corruption is applied at 5 different severity levels to the test sets of CIFAR10, CIFAR100 [22], and ImageNet [8], acting as source test sets, and allowing us to systematically evaluate the model’s performance under increasing degrees of image degradation.
@@
where $\text{sim}(\cdot)$ is the cosine similarity and $\tau$ is the softmax temperature from CLIP’s pre-training stage. The text features $\mathcal{Z}$, in zero-shot evaluation, are always pre-computed. We also draw a comparison with CLIP performance on respective source test sets and follow this implementation  [^1] https://github.com/LAION-AI/CLIP_benchmark. Throughout this paper, the same TTA problem setup is used, unless mentioned otherwise. This TTA setup differs from the TTA setups of [36, 15, 34].
Figure 2: Task-wise mean accuracy (%) of zero-shot CLIP across corruption severity levels. [Top]: ResNet-101 backbone. [Bottom]: ViT-B/16 backbone. Dashed lines show zero-shot CLIP performance on corresponding source test sets (clean).
@@ 3.1 Sensitivity of CLIP to image corruption severity
We analyze CLIP’s zero-shot performance by progressively increasing corruption severity for RN101 and ViT-B/16 vision backbones. We illustrate mean accuracy and overall performance in Figure 2. Despite CLIP’s robust multimodal feature space, accuracy drops significantly with an increase in corruption severity, regardless of the backbone. Specifically, for CIFAR-10C with a ViT-B/16 backbone, we observe accuracy as low as 37.92%, at a severity level of 5, for Gaussian noise. Similarly, for CIFAR-100C and ImageNet-C, the mean accuracy rates are as low as 35.79% and 24.51% at a severity level of 5, respectively. Results on additional vision backbones are in the Supplementary.
@@
(a) CIFAR-10C
(b) CIFAR-100C
(c) ImageNet-C
Figure 3: Mean classification accuracy (in %) across all the corruption types vs. accuracy on corresponding source test sets, evaluated using various prompt templates and across vision backbones.
@@ 3.2 Sensitivity of CLIP to prompt templates
In this analysis, we evaluate the impact of prompt engineering by providing “relevant” prompt templates to $f_{txt}$ for TTA. Each prompt adds context to help CLIP extract more relevant text features. We report the mean classification accuracy (in %) across RN101, ViT-B/16, ViT-B/32, and ViT-L/14 backbones at an image corruption severity level of 5 for all datasets, with results presented in Figure 3. For each dataset, the x-axis is the source accuracy and y-axis is the mean accuracy, across 15 tasks.
@@
As expected, RN101 performs worse than the ViT-based backbones, primarily due to its lack of global attention-based modeling inherent to transformers [50]. Therefore, for the remainder of the experiments in this paper, we focus on ViT-based backbones, specifically ViT-B/16.
Unsuitability of prompt template selection at test-time. Additionally, at test-time, it is impractical to perform prompt engineering or optimize prompt vectors since 1) Choosing different prompt templates for generating text features is extremely tedious and time-consuming. As discussed in Section 2, CLIP TTA works [36, 15, 34] use various prompt templates, making it unrealistic. 2) In real-time deployment involving prompt-tuning, such prompts cannot quickly estimate the distribution of incoming test batches. TPT [45] optimizes pre-trained text prompts for each test image, turning out to be suboptimal since the prompts are optimized ignorant of the distribution of the test image.
4 Proposed Methodology
Figure 4: Illustration of our proposed approach: BATCLIP not only adapts the visual encoder for highly discriminative image features but also promotes a strong alignment between image and text features by adapting the text encoder too, leading to improved performance following test-time adaptation. We adapt only the LayerNorm parameters of CLIP encoders.
The comprehensive analysis in Section 3 reveals that the zero-shot vision encoder $f_{vis}$ of CLIP is very sensitive to image corruption with increasing severity. Similarly, the performance of the text encoder $f_{txt}$ is invariant to the different text prompt templates and is also impractical to tune at test-time. Therefore, for effective adaptation, both the image and text encoders of CLIP need to be adapted to the incoming domain of test batches. We illustrate our proposed framework, BATCLIP, in Figure 4.